Chapter 25, Problem 004 pates of a spherical capacitor have radii 39.7 mm and 42 . must be the plate area of a parallel-plate capacitor with the same plate separation and capacitance? (a) Number- Zu (b) Number Units Units
Answer #1

a-spherical capacitance=

=1.1*10-10 *1.9*10-6

= 2.09*10-16 c

b- plane plate seperation= 3mm

capacitance =2.09*10-16 c

and we know capacitance =

2.09*10-16 = 8.8*10-12 * a /3*10-3

a= 0.7125*10-7m2

  a=7.125*10-8 m2